OpenStudy (anonymous):

On a die numbered 1 to 6, What is the probability of rolling a 2 three times in a row. 1/6 1/2 1/72 1/216

Directrix (directrix):

There are 6 outcomes when you toss a single die: 1, 2, 3, 4, 5, and 6 Of the six outcomes, 2 is one of them. The probability of tossing a "2" on ONE toss of a die is 1/6.

Directrix (directrix):

You are tossing the die THREE times hoping to get a 2 each time. Any idea on how to get that probability @undeadking1357

Directrix (directrix):

No.

OpenStudy (anonymous):

hmm I don't understand @Directrix

Directrix (directrix):

Each probability is 1/6. Hint: Probability Multiplication Principle.

Directrix (directrix):

First Toss has probability of 1/6 Second Toss has probability of 1/6 and Third Toss has probability of 1/6 So, P(2,2,2) = 1/6 * 1/6 * ? = ?

OpenStudy (anonymous):

oh the answer is 1/216 @Directrix

Directrix (directrix):

Correct.